Strontium chloride and sodium fluoride react to form strontium fluoride and sodium chloride, according to the reaction shown.


SrCl2(aq)+2NaF(aq)⟶SrF2(s)+2NaCl(aq)


What volume of a 0.190 M NaF solution is required to react completely with 417 mL

 of a 0.400 M 0.400 M SrCl2 solution?

How many moles of SrF2 are formed from this reaction?


1
Expert's answer
2022-04-07T03:55:04-0400

SrCl2(aq)+2NaF(aq)⟶SrF2(s)+2NaCl(aq)

CM = n / V

n (SrCl2) = 1/2 x n (NaF)

2 x (CM (NaF) x V (NaF)) = CM (SrCl2) x V (SrCl2)

2 x (0.19 x V (NaF)) = 0.400 x 0.417

V (NaF) = (0.400 x 0.417) / (0.19 x 2) = 0.44 L

n (SrF2) = n (SrCl2) = 0.400 x 0.417 = 0.17 mol